# Helios: An extremely low power event-based gesture recognition for always-on smart eyewear

### v1

@@ 5.1 Event Representation
The window size $L$ is important to choose properly: if it is too long, the TS can be overridden when a new event with the same polarity occurs at the same pixel location; if it is too short, it can be computationally expensive. For our experiments, we experimented with the ideal window size and used a fixed time length window for all micro-gestures.
@@ 5.2 Machine Learning Model
In this section we provide details on our machine learning model for the task of micro-gesture detection.
@@ 5.2.1 Model Interface
The input to our model is a TS as detailed in Section 5.1. The model predicts probabilities of the input TS containing each of the seven output classes, where the seven output classes contain both micro-gestures and some adversarial classes included to improve micro-gesture classification when testing on a real device.
@@ 5.2.2 Architecture Details
Our model, as shown in Figure 5, is a two stage model. The first stage takes TS as an input and predicts a bounding box of the hand in the input and a class prediction of whether there is a hand or not in the bounding box. The TS is then cropped to the bounding box. The second stage takes as input the cropped TS and the bounding box class prediction. It then predicts the probability of the cropped TS containing each of the micro-gestures. The final model output is then calculated by multiplying the micro-gesture probabilities with the probability that there was a hand in the TS and concatenating this with the probability of there not being a hand.
@@ 5.2.3 Loss Function
When training our model we use two loss functions. The first is a bounding box loss $\mathcal{L}_{\text{bbox}}$, which computes the mean squared error between predicted and true bounding boxes when the true bounding box has a hand in it and provides no signal when the true bounding box does not have a hand. The second loss function $\mathcal{L}_{\text{gesture}}$ is a sparse categorical cross entropy loss between the predicted and true micro-gestures classes. We found that $\mathcal{L}_{\text{bbox}}$ is required for two reasons. First, the signal from $\mathcal{L}_{\text{gesture}}$ is not strong enough to guide the model to predict bounding boxes that produce TS crops that have the hand centered. Secondly, without $\mathcal{L}_{\text{bbox}}$, the model first learns to distinguish the no hand class from the others and then begins to start learning to distinguish the other micro-gesture classes, which causes a drop in performance of the final model suggesting it got stuck in a local minimum.
@@ 5.2.5 Model Inference
At inference time the model requires an aggregated TS as the input. To avoid sampling issues where the micro-gesture occurred during a period that intersects two models inputs, we create an $L$ window-sized aggregated TS after every 80ms period. The intuition for this is that the model will be unsure on predictions where non-complete micro-gestures have occurred in the input frame, which will be assigned low probabilities by the model. On the other hand, the predicted probability will be higher when the micro-gesture occurred within the aggregation window. We overcome the issue of the model predicting a micro-gesture with low confidence by introducing a probability threshold at which a micro-gesture is deemed to have occurred.
@@ 5.3 Dataset collection for training ML gesture recognition system
Due to the lack of event stream datasets for hand pose estimation or micro-gesture detection and the difficulty of obtaining accurate ground-truth annotations on real data, we build a highly efficient event stream simulator to generate a large-scale synthetic events dataset with annotations. Our seven classes of interest include: double-pinch, left-swipe, right-swipe, pinch, resting hands, idle and random hand motion. We generated examples of simulated event data per micro-gesture class for training and validation. The exact number of samples for training and validation sets per gesture class is given in Figure 9. For our testing setup, we use live user testing as described in Section 6.4.
We build on the ESIM simulator [19] to generate training data for the gesture classes. Simulating an event camera requires access to a continuous representation of the visual signal at every pixel, which is not feasible in practice. To address this, previous works have proposed sampling the visual signal synchronously at a very high frame rate and using linear interpolation between samples to create a piecewise linear approximation of the continuous visual signal. We adopt this approach for simulating events, by sampling the visual signal from recorded videos of the micro-gesture being performed. An example of TS generated from simulated data for left swipe and right swipe micro-gestures is given in Figure 6.
@@ 6.4.1 Precision and Recall
We evaluate the performance of our ML model with live user testing on 20 participants. For each participant, we collected the results for double-pinch, left micro-gesture and right micro-gestures. Each gesture is performed in a randomized order by a user 10 times, thereby leading to 30 gestures per user. The time between each gesture is 1.5s and the time given to complete a gesture is 2s. The precision and recall for 20 users and the learning curve for 5 naive users is given in Table 2. Our recall scores range from 54% to 86%, and precision scores range between 74% to 92%. The reason for double pinch scores to be lower than the swipes are that in the test app we consider the pinch to be distinct from a double pinch, and we do not optimize the TS to be able to cleanly distinguish between them. We expect the double pinch performance to improve if we set the pinch classification threshold to be very high.

### v2

@@ 5.1 Event Representation
The window size $L$ is important to choose properly: if it is too long, the TS can be overridden when a new event with the same polarity occurs at the same pixel location; if it is too short, it can be computationally expensive. For our experiments, we experimented with the ideal window size and used a fixed time length window for all microgestures.
@@ 5.2 Machine Learning Model
In this section we provide details on our machine learning model for the task of microgesture detection.
@@ 5.2.1 Model Interface
The input to our model is a TS as detailed in Section 5.1. The model predicts probabilities of the input TS containing each of the seven output classes, where the seven output classes contain both microgestures and some adversarial classes included to improve microgesture classification when testing on a real device.
@@ 5.2.2 Architecture Details
Our model, as shown in Figure 5, is a two stage model. The first stage takes TS as an input and predicts a bounding box of the hand in the input and a class prediction of whether there is a hand or not in the bounding box. The TS is then cropped to the bounding box. The second stage takes as input the cropped TS and the bounding box class prediction. It then predicts the probability of the cropped TS containing each of the microgestures. The final model output is then calculated by multiplying the microgesture probabilities with the probability that there was a hand in the TS and concatenating this with the probability of there not being a hand.
@@ 5.2.3 Loss Function
When training our model we use two loss functions. The first is a bounding box loss $\mathcal{L}_{\text{bbox}}$, which computes the mean squared error between predicted and true bounding boxes when the true bounding box has a hand in it and provides no signal when the true bounding box does not have a hand. The second loss function $\mathcal{L}_{\text{gesture}}$ is a sparse categorical cross entropy loss between the predicted and true microgestures classes. We found that $\mathcal{L}_{\text{bbox}}$ is required for two reasons. First, the signal from $\mathcal{L}_{\text{gesture}}$ is not strong enough to guide the model to predict bounding boxes that produce TS crops that have the hand centered. Secondly, without $\mathcal{L}_{\text{bbox}}$, the model first learns to distinguish the no hand class from the others and then begins to start learning to distinguish the other microgesture classes, which causes a drop in performance of the final model suggesting it got stuck in a local minimum.
@@ 5.2.5 Model Inference
At inference time the model requires an aggregated TS as the input. To avoid sampling issues where the microgesture occurred during a period that intersects two models inputs, we create an $L$ window-sized aggregated TS after every 80ms period. The intuition for this is that the model will be unsure on predictions where non-complete microgestures have occurred in the input frame, which will be assigned low probabilities by the model. On the other hand, the predicted probability will be higher when the microgesture occurred within the aggregation window. We overcome the issue of the model predicting a microgesture with low confidence by introducing a probability threshold at which a microgesture is deemed to have occurred.
@@ 5.3 Dataset collection for training ML gesture recognition system
Due to the lack of event stream datasets for hand pose estimation or microgesture detection and the difficulty of obtaining accurate ground-truth annotations on real data, we build a highly efficient event stream simulator to generate a large-scale synthetic events dataset with annotations. Our seven classes of interest include: double-pinch, left-swipe, right-swipe, pinch, resting hands, idle and random hand motion. We generated examples of simulated event data per microgesture class for training and validation. The exact number of samples for training and validation sets per gesture class is given in Figure 9. For our testing setup, we use live user testing as described in Section 6.4.
We build on the ESIM simulator [19] to generate training data for the gesture classes. Simulating an event camera requires access to a continuous representation of the visual signal at every pixel, which is not feasible in practice. To address this, previous works have proposed sampling the visual signal synchronously at a very high frame rate and using linear interpolation between samples to create a piecewise linear approximation of the continuous visual signal. We adopt this approach for simulating events, by sampling the visual signal from recorded videos of the microgesture being performed. An example of TS generated from simulated data for left swipe and right swipe microgestures is given in Figure 6.
@@ 6.4.1 Precision and Recall
We evaluate the performance of our ML model with live user testing on 20 participants. For each participant, we collected the results for double-pinch, left microgesture and right microgestures. Each gesture is performed in a randomized order by a user 10 times, thereby leading to 30 gestures per user. The time between each gesture is 1.5s and the time given to complete a gesture is 2s. The precision and recall for 20 users and the learning curve for 5 naive users is given in Table 2. Our recall scores range from 54% to 86%, and precision scores range between 74% to 92%. The reason for double pinch scores to be lower than the swipes are that in the test app we consider the pinch to be distinct from a double pinch, and we do not optimize the TS to be able to cleanly distinguish between them. We expect the double pinch performance to improve if we set the pinch classification threshold to be very high.